Write a balanced equation for the overall cell reaction and calculate the standard cell potential, E^0cell.




A. Half-reactions: Cd^2+(aq) + 2e^- —> Cd(s) Cr^2+(aq) + 2e^- —> Cr(s)




Cell reaction:




E^0 cell





B. Half-reactions: Fe^2+(aq) + 2e^- —> Fe(s) Br2(g) + 2e^- —> 2Br^-(aq)




Cell reaction:




E^0cell

1
Expert's answer
2022-06-13T17:02:13-0400

The overall cell potential can be calculated by using the equation E0cell=E0red−E0oxid.



A.)

CdSO4 + Cr = CrSO4 + Cd

E0 cell = -0,74 - (-0,4) = -0,34 V


B.)

Fe + Br2 = FeBr2

E0 cell = -0,44 - 1,078 = -1,518 V